Grade 9General Physics

If the initial velocity is 20m/sec ,final velocity is 0m/sec,and the distance of 10cm.what will be the retardation?

Kalpana
9 Years agoGrade 9

1 Answer

hero
9 Years ago
initial velocity(u)=20m/s
final velocity(v)=0m/s
distance(s)=10m
v2-u2=2as
02-202=2×a×10
On solving
a=-20m/s2
Hope it helps